OpenStudy (anonymous):

given that x=ρsinϕcosθ,y=ρsinϕsinθ,z=ρcosϕ prove (x_phi_^2)+(y_phi_^2)+(z_phi_^2)=ρ^2

OpenStudy (anonymous):

Do each at a time. \[x^2 = (\rho*sin(\phi)*cos(\theta))^2=\rho^2*sin^2(\phi)*cos^2(\theta)\] \[y^2 = (\rho*sin(\phi)*sin(\theta))^2=\rho^2*sin^2(\phi)*sin^2(\theta)\] \[z^2 = (\rho*cos(\phi))^2=\rho^2*cos^2(\phi)\] Then sum them up \[\rho^2*sin^2(\phi)*cos^2(\theta) +\rho^2*sin^2(\phi)*sin^2(\theta)+ \rho^2*cos^2(\phi)\] Notice you can pull out a rho^2 and group things up. \[\rho^2*[sin^2(\phi)*cos^2(\theta) +sin^2(\phi)*sin^2(\theta)+ cos^2(\phi)]\]

OpenStudy (anonymous):

Then you can group the \[sin^2(\phi)\] and you get the identity \[cos^2(\theta) +sin^2(\theta) =1\] \[\rho^2*[sin^2(\phi)*(cos^2(\theta) +sin^2(\theta))+ cos^2(\phi)]\] simplify that you get another identity of one \[\rho^2*[sin^2(\phi)+ cos^2(\phi)]\]

OpenStudy (anonymous):

Wow, thank you. That helped me a lot. I kept trying to find a way to make partial derivatives to work. Glad to know I was just over thinking it. =]